# Premerger phenomena in neutron-star binary coalescences

### v2

@@ VI.4.3 Ocean modes
Ocean modes refer rather literally to waves (as for surface waves on Earth’s oceans) that occur in the thin, penultimate layer of the star (before atmosphere), which is typically a few meters to tens of meters thick [262, 371]. Given the low density of the region, these modes are especially affected by magnetic field, temperature gradients, gravitational forces, and most importantly composition. Ocean modes are important in the context of X-ray bursts, as the thermonuclear reactions that heat up the ocean layer can also instigate drifts that modulate the X-ray emission [see Table 4 and references in 426]; such drifting can be used to study ocean and atmosphere structure [411, 81, 285].
Ocean modes tend to have frequencies of $\sim$ tens of Hz at most, and thus could be excited at early times in a merger. Similar to the $g$-mode case, there is a competition between the excitation window (which is long for low frequencies) and overlap (which tends to be small at low frequencies). This was investigated by Sullivan et al. 373, who found largest overlaps for carbon atmospheres with frequencies $\sim 16$ Hz. These authors also found that energy deposits could reach $\lesssim 10^{47}$ erg over the inspiral duration, decreasing by a few orders of magnitude if elasticity or a crust with heavier nuclei (e.g. oxygen or iron) is considered. Atmospheres composed of lighter elements may be expected in a system with a history of accretion; “rapid” rotation may thus coincide with larger ocean impact (Sec. II.2). Again like the $g$-mode case, there should also be a contest between mode frequencies and the local reaction rate, though the reaction rate in the ocean can be quite different from that in the crust or core. For heavier atmospheres with sub-Hz mode frequencies, space-based GW interferometers may be especially useful since such resonances could occur long before the LVK window and possibly coincide with multimessenger activity [373]. Whether such modes could ever instigate crustal failure is unclear, though emissions may still result from particular accelerations via tidal-wave motions [372].
Aside from the crust-core interface, there is also a crust-ocean interface that can be considered (the $i_{1}$ mode). These can potentially grow to large amplitudes; as noted by Sullivan et al. 372, while crustal failure may be difficult to achieve through an initial resonance as $i_{1}$ overlaps are not so large [see Figure 9 in 302], crustal inhomogeneities may arise from such local failures [see also 207]. This may prime the crust for failures down the line, through either repeated resonances of the same mode, in an eccentric or dynamically-evolving system, or higher-frequency (e.g. $g$-) modes.
@@
The f-undamental fluid oscillation mode is a member of the acoustic ($p$-) family, in which the main restoring force is the hydrostatic pressure. This mode has no nodes, and thus couples very strongly to the tidal field which also has no nodes, as noted by Thorne 402 and others. The linear eigenfrequencies of these modes typically range from 1 kHz to several kHz, depending on the mass, radius, and internal structure of the star. However, several studies have noted an insensitivity to the EOS [30, 31], allowing for universal relations to be written down. Such aspects are powerful in the sense that measuring an $f$-mode frequency would give valuable insight onto the compactness of the system in an EOS-independent way. For a complete list of such relations, including the effects of rotation, we refer the reader to Krüger et al. 225, Manoharan and Kokkotas 257.
@@
Inertial (i.e. $r$-) modes are quasi-axial modes, degenerating in the static limit, with relatively weak coupling to the tidal field. However, in systems that are rotating very rapidly ($\nu\gg 100$ Hz), the hybrid character allows the modes to exert non-negligible strain on the crust [238, 383], though it is expected that such systems are rare [cf. Sec. II.2 449]. They are generically unstable to rotational instabilities [23, 145, 32, 219] because their co-rotating frame frequencies are less than $2\Omega_{s}$ (see equation 24) and in this way they may be important in limiting the maximum rotation of neutron stars through radiation-reaction [29]. They can be potentially excited in neutron star glitches [37], the predicted amplitudes of which may be detectable with LVK given $\gtrsim$ months of folding for Galactic sources [e.g. PSR J0537-691; 26]. Mathematical considerations related to tidal and magnetic corrections to the inertial spectrum are considered in Suvorov and Kokkotas 383, and are similarly not expected to be important unless magnetars are present in the merger (cf. Secs. II.3 and IV.2). In principle, $r$-modes could also incite dynamo activity prior to merger if reaching a large enough amplitude [335, see also Sec. VI.5].
@@ VI.5 Late-stage dynamos?
Tidal resonances, aside from draining energy from the orbit, also add angular momentum to the stellar interior (Sec. IV.3.1). The resulting velocity pattern is tied to the mode eigenfunction and generally implies differentially rotating cavities, $\Omega=\Omega(t,\boldsymbol{x})$. One such example, computed self-consistently by Suvorov et al. 386, is shown in Figure 21 for two different $g$-mode resonances. The snapshots are taken at the moment of peak positive amplitude in the crust. Generally speaking, low-overtone $g$-modes have a monotonic radial profile in the crust, leading to either $\partial_{r}\Omega<0$ or $\partial_{r}\Omega>0$ everywhere there, depending on mode phase.
Figure 21: Crustal angular velocities at fixed spherical angles, $\Omega(r)$, for $g_{1}$ (black, solid) and $g_{2}$ (blue, dashed) modes, at a time slice corresponding to the maximum resonant amplitude. The star has an initial angular velocity of $\Omega_{s}=200$ rad/s with $M=1.6M_{\odot}$, and $R=11.26\,$km with a hybrid APR4 + Douchin and Haensel 127 (for the crust) EOS. From Suvorov et al. 386.
@@
Fields of strength depicted in Fig. 22 would allow for the magnetic energy budget to match the observed luminosities of precursors, possibly resolving the mystery of why precursors are rare: fast stars are needed. Suvorov et al. 386 estimated that the conditions necessary for the MRI are satisfied in $\sim 5\%$ of merger events (see Appendix A therein). It may therefore be that magnetar-level fields do not persist into merger, but rather the field is generated. While appealing, some issues remain with this interpretation however. The main one being that it is difficult to imagine a coherent field being produced; rather one may expect a turbulent, highly-tangled structure to emerge, as is more typical from MRI simulations [e.g. 212]. The process of extracting magnetoelastic energy is thus more complicated, and further work is needed to address the topological development of the field. A second issue is that estimates thus far have been carried out in (GR)MHD, but really elastic terms are necessary to adequately describe the crustal system. Whether it is possible for the MRI to activate in an elastic cavity, rather than fluid interior, is not obvious. Efforts in this direction are underway. Other dynamos would worth investigating, like those with a chiral flavor [110] triggered by tide-related processes that force the system out of thermodynamic equilibrium prior to merger [176].
Figure 22: Dynamo amplifications from either a local MHD analysis (solid curves) or setting equipartition between magnetic and shear energies (dashed curves; upper limits), as triggered by either $g_{1}$- or $g_{2}$-modes in a star with $\Omega_{s,0}=200$ rad/s. The same macro- and microparameters used in Fig. 21 are used. From Suvorov et al. 386.
@@ VI.6 Post-merger models
If the main GRB episode is produced by (the standard mechanism involving) synchrotron radiation in an internal shock or magnetic dissipation zone, a (thermal) precursor may be expected either as the shock breaks out from the surrounding ejecta or as some (neutrino-anti-neutrino) fireball reaches the photosphere [271, 424, 422]. Alternatively, it may be that an incipient jet fails to break out (or “chokes”) which can produce some GRB-like emission prior to the main event, where a renewed jet successfully escapes. In the latter case, a range of spectra are plausible depending on jet nature. A newborn magnetar may also launch a powerful wind independently from the jet, which may interact with the surroundings and produce X- and gamma-ray activity [e.g. 103]. Overall this Review is however concerned with premerger phenomena, and thus we do not discuss these possibilities further. The interested reader is directed to Ciolfi 91, Burns 74, Sarin and Lasky 345 and references therein.
@@ VII Multimessenger outlook
The previous three sections have been devoted to understanding gravitational and gamma-ray emissions in premerger systems. These can be studied through tidal and magnetic interaction theory, with a careful of the stellar micro- and macrophysics. Here, by contrast, we turn briefly to other multimessenger elements, expected in the radio and X-ray bands, together with neutrino counterparts. We intentionally keep this section rather short: such elements have been described elsewhere [see, for instance 252] and our main focus is on the gamma- and GW aspects.
@@
    Gamma-ray precursors. Depending on the ignition mechanism, different kinds of information be may discernible. For premerger precursors, we have argued the resonant failure picture can adequately explain all the observational characteristics (Sec. VI.3), though admittedly this is due to the huge range of QNM properties that neutron stars can exhibit. As resonances are obviously tied to the mode spectra, everything above applies here too but in the gamma-ray band. Similarly, since the activity of these modes in the crust is the relevant aspect here, microphysical inputs become critical. With the above three (Love number, dynamical tides, and precursors) one may thus learn about both macro- and microphysical elements of the stars taking part in a merger. This can be combined self-consistently with the properties of the post-merger remnant to deduce generative elements of GRBs (e.g. heavier stars will more likely promptly collapse leading to faster jet break-out; see Zhang 446 and Tab. 4).
  4. 4.
@@
    Radio flares. Radio activity can be incited premerger from a few different channels. For instance, there may be shock-powered radio emissions through interactions taking place in the accelerating, binary wind left in the wake of the inspiral [270, 363]. These are likely to be in the form of FRBs or “giant pulse” like phenomena in the $\sim$ GHz band. Another possibility discussed by Cooper et al. 95 is that acceleration zones may form in regions of interwoven magnetic fields, which could produce coherent, millisecond bursts in radio frequencies that are theoretically observable out to $\sim$ Gpc distances. Such observations could reveal information about the magnetospheric plasma and radio activation mechanisms, which can be used to deduce information about the general pulsar engine and all that can convey about stellar structure [see, e.g. 58]. There is also the possibility of postmerger radio activity associated with a neutron-star collapse, as described in Sec. VI.1. The Square Kilometer Array (SKA) will go a long way towards detection prospects.
  5. 5.
@@
    X-ray flares. We have earlier described how the emission mechanism for GRB precursors resembles that of giant flares from magnetars in terms of energy extraction and propagation. Given that X-ray activity from magnetars is commonplace (and arguably a defining feature), the same applies for premerger stars with strong fields. X-ray emission mechanisms in a premerger system are described by Beloborodov 51, the key ingredient being the nonlinear development of magnetospheric Alfvén waves [see also 278]. Such observations may reveal crucial information about the local magnetic field strength and radiation transport physics, from which information about neutron-star evolutionary pathways can be deduced. X-ray activity may also be related to magnetic instabilities arising by late stage dynamo activity or magnetic reconfiguration more generally [see Sec. VI.5 and also 259, 377, for instance]. The planned ECLAIRs telescope will help enable searches for X-rays out to cosmological distances.
  6. 6.
@@
    Neutrinos. Neutrinos could be emitted from a premerger system through at least two distinct means. One involves tidal heating. As described in Sec. IV.5, rapid episodes of heating could lead to chemical imbalance and the production of neutrinos which could be theoretically observed for a close merger. Such observations would provide important information on the microphysical heat capacity. The second mechanism is indirectly via the production of cosmic rays. As discussed by Coppin and van Eijndhoven 96, it is thought that GRBs could be sources of ultra-high-energy cosmic rays. In this case, the interaction of high-energy protons in relativistic wind wakes (as above) near the source could produce neutrinos carrying a non-negligible fraction of the proton energy. Ice-Cube has reported detection of some cosmic neutrinos [1, 2, 5], while the sources are only identified in a few cases; next-generation neutrino observatories (like IceCube-Gen2; Aartsen et al. 3) will synergise with other instruments to better constrain the sources [e.g. 282]. In addition, such observations could be used to place constraints on binary neutron-star merger abundances and provide tests of the standard model [22]. For instance, neutrinos are thought to carry magnetic moments [148, 358] through which they interact with the ambient magnetic field that depends on whether they are of a Dirac or Majorana nature [160]. Thus, the neutrino flux from a strongly magnetised environment, such as the engine of GRBs [427, 417], may encode neutrino’s nature [e.g. 70]. An excellent review of future neutrino observatories, and instruments capable of follow-up, are described in Tables 1 and 2 of Guépin et al. 172.
@@ VII.1 A brief look at future possibilities
One of the key ways in which multimessenger astrophysics will be propelled forward in the future comes from next-generation GW interferometers. Although there are a number of proposed technologies, some of the most notable are (i) the Einstein Telescope [ET; 254, 322, 184, 68]. The sensitivity band of ET is similar to that of the LVK network, though notably deeply with considerable improvement in the $\lesssim 10$ Hz range. The reason is that the gravitational gradient (aka “Newtonian”), seismic, and thermal noises are less of a problem for ET, as it will be built underground and include cryogenic technologies to reduce thermal vibrations [322]. This latter band is relevant to capture the onset of low-frequency oscillations and track the orbit for longer periods of time. (ii) Cosmic Explorer [CE; 6, 334, 133]. This detector is similar to ET in terms of sensitivity band, though is planned to go even deeper. CE may detect thousands of neutron stars in merger [134], and thus lead to a true era where statistical methods can nail down the neutron-star EOS. (iii) The Laser Interferometer Space Antenna (LISA). LISA is a planned space-based detector that could capture the very early aspects of inspiral while the orbit is at $\gtrsim$ mHz frequencies [21]. Wide binary elements could be useful for ocean resonances [373, 372], low frequency $g$-modes (assuming these are not washed out by reaction rates; Sec. III.3) and the testing of spacetime structure generally [see, e.g. 115, 116].
@@
With CE and ET, many of the high-amplitude QNMs of a merger remnant may be resolvable out to cosmological distances. Aside from providing tests of GR directly [see, e.g. 218, 57, 220, 392], this could be connected with GRB observations to deduce remnant nature and its impact on jet structure; see Chirenti et al. 88, who reported the discovery of $\sim$ kHz modulations in the short GRBs 910711 and 931101B. Identifying oscillation frequencies in both the remnant and the premerger stars (via dynamical tides) would prove a powerful probe of fundamental physics. In a similar vein, GWs from a resonant failure itself may be detectable with a sufficiently sensitive telescope for a Galactic binary merger [451]. This would allow for the precursor-merger waiting time to be cleanly isolated, removing uncertainties relating to jet-break out (see Sec. VI.1).
@@
Innovations on the electromagnetic side are no less promising. For instance, with the development of the SKA and next-generation X-ray observatories the types of premerger precursors could increase significantly. Although we have limited our use of the word precursor to specifically mean the gamma-ray flashes occurring before the main event (see Tab. 3), it is hoped that such terminology will prove ambiguous in the future. Tables 1 and 2 of Guépin et al. 172 provide a summary of next-generation telescopes that could coincidentally detect signals associated with neutrino bursts originating from a binary merger. One intriguing possibility comes from the planned POLAR-2 detector, which should able to measure polarization and similar effects related to the propagation of gamma-rays [a survey of other next-generation gamma-ray facilities are described in Section 2 of 195]. This could be used to probe the local magnetic field strength in the region where the rays were produced. This would allow one to test, for instance, the extent to which non-thermality is tied to magnetars in mergers (see also Sec. IV.4).

### v3

@@ VI.4.3 Ocean modes
Ocean mode oscillations in neutron stars refer rather literally to waves (similar to the surface waves on Earth’s oceans) that occur in the thin, penultimate layer of the star (before atmosphere), which is typically a few meters to tens of meters thick [272, 385]. Given the low density of the region, these modes are affected by factors such as the magnetic field, temperature gradients, gravitational forces, and most importantly composition. These modes are important in the context of X-ray bursts, as the thermonuclear reactions that heat up the ocean layer can also instigate drifts that modulate the X-ray emission [see Table 4 and references in 440]; such drifting can be used to study ocean and atmosphere structure [425, 80, 297].
Ocean modes tend to have frequencies of $\sim$ tens of Hz at most, and thus could be excited at early times in a merger (see Fig. 1). Similar to the $g$-mode case, there is a competition between the excitation window (which is long for low frequencies) and overlap (which tends to be small at low frequencies). This was investigated by Sullivan et al. 387, who found largest overlaps for carbon atmospheres with frequencies $\sim 16$ Hz. These authors also found that energy deposits could reach $\lesssim 10^{47}$ erg over the inspiral duration, decreasing by a few orders of magnitude if elasticity or a crust with heavier nuclei (e.g. oxygen or iron) is considered. Atmospheres composed of lighter elements may be expected in a system with a history of accretion; one may therefore expect “rapid” rotation to coincide with larger ocean impact (Sec. II.2). Again like the $g$-mode case, there should also be a contest between mode frequencies and the local reaction rate, though the reaction rate in the ocean can be quite different from that in the crust or core. For heavier atmospheres with sub-Hz mode frequencies, space-based GW interferometers may be especially useful since such resonances could occur long before the LVK window and possibly coincide with multimessenger activity [387]. Whether such modes could ever instigate crustal failure is unclear, though emissions may still result from particular accelerations via tidal-wave motions [386].
Aside from the crust-core interface, there is also a crust-ocean interface that can be considered (the $i_{1}$ mode). These can potentially grow to large amplitudes; as noted by Sullivan et al. 386, while crustal failure may be difficult to achieve through an initial resonance as $i_{1}$ overlaps are not so large [see Figure 9 in 315], crustal inhomogeneities may arise from such local failures [see also 214]. This may prime the crust for failures down the line, through either repeated resonances of the same mode, in an eccentric or dynamically-evolving system, or higher-frequency (e.g. $g$-) modes.
@@
The f-undamental fluid oscillation mode is a member of the the acoustic ($p$-) family, in which the main restoring force is the hydrostatic pressure. This mode has no nodes, and thus couples very strongly to the tidal field which also has no nodes, as noted by Thorne 416 and others before. The linear eigenfrequencies of these modes typically range from 1 kHz to several kHz, depending on the mass, radius, and internal structure of the neutron star. However, several studies have noted an insensitivity to the EOS [30, 31], allowing for universal relations to be written down. Such aspects are powerful in the sense that measuring an $f$-mode frequency would give valuable insight onto the compactness of the system in an EOS-independent way. For a complete list of such relations, including the effects of rotation, we refer the reader to Krüger et al. 234, Manoharan and Kokkotas 266.
@@
Inertial (i.e. $r$-) modes are quasi-axial modes, degenerating in the static limit, with relatively weak coupling to the tidal field. However, in systems that are rotating very rapidly ($\nu\gg 100$ Hz), the hybrid character allows the modes to exert non-negligible strain on the crust [247, 397], though it is expected that such systems are rare [cf. Sec. II.2 464]. They are generically unstable to rotational instabilities [23, 148, 32, 227] because their co-rotating frame frequencies are less than $2\Omega_{s}$ (see equation 24) and in this way they may be important in limiting the maximum rotation of neutron stars through radiation-reaction [29]. They can be potentially excited in neutron star glitches [37], the predicted amplitudes of which may be detectable with LVK given $\gtrsim$ months of folding for Galactic sources [e.g. PSR J0537-691; 26]. Mathematical considerations related to tidal and magnetic corrections to the inertial spectrum are considered in Suvorov and Kokkotas 397, and are similarly not expected to be important unless magnetars are present in the merger (cf. Secs. II.3 and IV.2). In principle, $r$-modes could also incite dynamo activity prior to merger if reaching a large enough amplitude [349, see also Sec. VI.5].
@@ VI.5 Late-stage dynamos?
Tidal resonances, aside from draining energy from the orbit, also add angular momentum to the stellar interior (Sec. IV.3.1). The resulting velocity pattern is tied to the mode eigenfunction and generally implies differentially rotating cavities, $\Omega=\Omega(t,\boldsymbol{x})$. One such example, computed self-consistently by Suvorov et al. 400, is shown in Figure 21 for two different $g$-mode resonances. The snapshots are taken at the moment of peak positive amplitude in the crust. Generally speaking, low-overtone $g$-modes have a monotonic radial profile in the crust, leading to either $\partial_{r}\Omega<0$ or $\partial_{r}\Omega>0$ everywhere there, depending on mode phase.
Figure 21: Crustal angular velocities at fixed spherical angles, $\Omega(r)$, for $g_{1}$ (black, solid) and $g_{2}$ (blue, dashed) modes, at a time slice corresponding to the maximum resonant amplitude. The star has an initial angular velocity of $\Omega_{s}=200$ rad/s with $M=1.6M_{\odot}$, and $R=11.26\,$km with a hybrid APR4 + Douchin and Haensel 128 (for the crust) EOS. From Suvorov et al. 400.
@@
Fields of strength depicted in Fig. 22 would allow for the magnetic energy budget to match the observed luminosities of precursors, possibly resolving the mystery of why precursors are rare: fast stars are needed. Suvorov et al. [400] estimated that the conditions necessary for the MRI are satisfied in $\sim 5\%$ of merger events (see Appendix A therein). It may therefore be that magnetar-level fields do not persist into merger, but rather the field is generated. While appealing, some issues remain with this interpretation however. The main one being that it is difficult to imagine a coherent field being produced; rather one may expect a turbulent, highly-tangled structure to emerge, as is more typical from MRI simulations [e.g. 219]. The process of extracting magnetoelastic energy is thus more complicated, and further work is needed to address the topological development of the field. A second issue is that estimates thus far have been carried out in (GR)MHD, but really elastic terms are necessary to adequately describe the crustal system. Whether it is possible for the MRI to activate in an elastic cavity, rather than fluid interior, is not obvious. Efforts in this direction are underway. Other dynamos would worth investigating, like those with a chiral flavor [110] triggered by tide-related processes that force the system out of thermodynamic equilibrium prior to merger [182].
Figure 22: Dynamo amplifications from either a local MHD analysis (solid curves) or setting equipartition between magnetic and shear energies (dashed curves; upper limits), as triggered by either $g_{1}$- or $g_{2}$-modes in a star with $\Omega_{s,0}=200$ rad/s. The same macro- and microparameters used in Fig. 21 are used. From Suvorov et al. 400.
@@ VI.6 Post-merger models
If the main GRB episode is produced by (the standard mechanism involving) synchrotron radiation in an internal shock or magnetic dissipation zone, a (thermal) precursor may be expected either as the shock breaks out from the surrounding ejecta or as some (neutrino-anti-neutrino) fireball reaches the photosphere [282, 438, 436]. Alternatively, it may be that an incipient jet fails to break out (or “chokes”) which can produce some GRB-like emission prior to the main event, where a renewed jet successfully escapes. In the latter case, a range of spectra are plausible depending on jet nature. A newborn magnetar may also launch a powerful wind independently from the jet, which may interact with the surroundings and produce X- and gamma-ray activity [e.g. 102]. Overall this Review is however concerned with premerger phenomena, and thus we do not discuss these possibilities further. The interested reader is directed to Ciolfi 90, Burns 73, Sarin and Lasky 359 and references therein.
@@ VII Multimessenger outlook
The previous three sections have been devoted to understanding gravitational and gamma-ray emissions in premerger systems. These can be studied through tidal and magnetic interaction theory, with a careful of the stellar micro- and macrophysics. Here, by contrast, we turn briefly to other multimessenger elements, expected in the radio and X-ray bands, together with neutrino counterparts. We intentionally keep this section rather short: such elements have been described elsewhere (see, for instance [261]) and our main focus is on the gamma- and GW aspects.
@@ 3. $\boldsymbol{\star}$
    Gamma-ray precursors. Depending on the ignition mechanism, different kinds of information be may discernible. For premerger precursors, we have argued the resonant failure picture can adequately explain all the observational characteristics (Sec. VI.3), though admittedly this is due to the huge range of QNM properties that neutron stars can exhibit. As resonances are obviously tied to the mode spectra, everything above applies here too but in the gamma-ray band. Similarly, since the activity of these modes in the crust is the relevant aspect here, microphysical inputs become critical. With the above three (Love number, dynamical tides, and precursors) one may thus learn about both macro- and microphysical elements of the stars taking part in a merger. This can be combined self-consistently with the properties of the post-merger remnant to deduce generative elements of GRBs (e.g. heavier stars will more likely promptly collapse leading to faster jet break-out; see Zhang 461 and Tab. 4).
  4. $\boldsymbol{\star}$
@@
    Radio flares. Radio activity can be incited premerger from a few different channels. For instance, there may be shock-powered radio emissions through interactions taking place in the accelerating, binary wind left in the wake of the inspiral [281, 377]. These are likely to be in the form of FRBs or “giant pulse” like phenomena in the $\sim$ GHz band. Another possibility discussed by Cooper et al. [94] is that acceleration zones may form in regions of interwoven magnetic fields, which could produce coherent, millisecond bursts in radio frequencies that are theoretically observable out to $\sim$ Gpc distances. Such observations could reveal information about the magnetospheric plasma and radio activation mechanisms, which can be used to deduce information about the general pulsar engine and all that can convey about stellar structure [see, e.g. 57]. There is also the possibility of postmerger radio activity associated with a neutron-star collapse, as described in Sec. VI.1. The Square Kilometer Array (SKA) will go a long way towards detection prospects.
  5. $\boldsymbol{\star}$
@@
    X-ray flares. We have earlier described how the emission mechanism for GRB precursors resembles that of giant flares from magnetars in terms of energy extraction and propagation. Given that X-ray activity from magnetars is commonplace (and arguably a defining feature), the same applies for premerger stars with strong fields. X-ray emission mechanisms in a premerger system are described by [50], the key ingredient being the nonlinear development of magnetospheric Alfvén waves [see also 290]. Such observations may reveal crucial information about the local magnetic field strength and radiation transport physics, from which information about neutron-star evolutionary pathways can be deduced. X-ray activity may also be related to magnetic instabilities arising by late stage dynamo activity or magnetic reconfiguration more generally (see Sec. VI.5 and also [269, 391] for instance). The planned ECLAIRs telescope will help enable searches for X-rays out to cosmological distances.
  6. $\boldsymbol{\star}$
@@
    Neutrinos. Neutrinos could be emitted from a premerger system through at least two distinct means. One involves tidal heating. As described in Sec. IV.5, rapid episodes of heating could lead to chemical imbalance and the production of neutrinos which could be theoretically observed for a close merger. Such observations would provide important information on the microphysical heat capacity. The second mechanism is indirectly via the production of cosmic rays. As discussed by Coppin and van Eijndhoven [95], it is thought that GRBs could be sources of ultra-high-energy cosmic rays. In this case, the interaction of high-energy protons in relativistic wind wakes (as above) near the source could produce neutrinos carrying a non-negligible fraction of the proton energy. Ice-Cube has reported detection of some cosmic neutrinos [1, 2, 5], while the sources are only identified in a few cases; next-generation neutrino observatories (like IceCube-Gen2; [3]) will synergise with other instruments to better constrain the sources [e.g. 294]. In addition, such observations could be used to place constraints on binary neutron-star merger abundances and provide tests of the standard model [22]. For instance, neutrinos are thought to carry magnetic moments [151, 372] through which they interact with the ambient magnetic field that depends on whether they are of a Dirac or Majorana nature [165]. Thus, the neutrino flux from a strongly magnetised environment, such as the engine of GRBs [441, 431], may encode neutrino’s nature [e.g. 69]. An excellent review of future neutrino observatories, and instruments capable of follow-up, are described in Tables 1 and 2 of Guépin et al. [177].
@@ VII.1 A brief look at future possibilities
One of the key ways in which multimessenger astrophysics will be propelled forward in the future comes from next-generation GW interferometers. Although there are a number of proposed technologies, some of the most notable are (i) the Einstein Telescope (ET; [263, 335, 190, 67]). The sensitivity band of ET is similar to that of the LVK network, though notably deeply with considerable improvement in the $\lesssim 10$ Hz range. The reason is that the gravitational gradient (aka “Newtonian”), seismic, and thermal noises are less of a problem for ET, as it will be built underground and include cryogenic technologies to reduce thermal vibrations [335]. This latter band is relevant to capture the onset of low-frequency oscillations and track the orbit for longer periods of time. (ii) Cosmic Explorer (CE; [6, 348, 135]). This detector is similar to ET in terms of sensitivity band, though is planned to go even deeper. CE may detect thousands of neutron stars in merger [136], and thus lead to a true era where statistical methods can nail down the neutron-star EOS. (iii) The Laser Interferometer Space Antenna (LISA). LISA is a planned space-based detector that could capture the very early aspects of inspiral while the orbit is at $\gtrsim$ mHz frequencies [21]. Wide binary elements could be useful for ocean resonances [387, 386], low frequency $g$-modes (assuming these are not washed out by reaction rates; Sec. III.3) and the testing of spacetime structure generally [see, e.g. 115, 116].
@@
With CE and ET, many of the high-amplitude QNMs of a merger remnant may be resolvable out to cosmological distances. Aside from providing tests of GR directly [see, e.g. 226, 56, 228, 406], this could be connected with GRB observations to deduce remnant nature and its impact on jet structure; see Chirenti et al. 87, who reported the discovery of $\sim$ kHz modulations in the short GRBs 910711 and 931101B. Identifying oscillation frequencies in both the remnant and the premerger stars (via dynamical tides) would prove a powerful probe of fundamental physics. In a similar vein, GWs from a resonant failure itself may be detectable with a sufficiently sensitive telescope for a Galactic binary merger [466]. This would allow for the precursor-merger waiting time to be cleanly isolated, removing uncertainties relating to jet-break out (see Sec. VI.1).
@@
Innovations on the electromagnetic side are no less promising. For instance, with the development of the SKA and next-generation X-ray observatories the types of premerger precursors could increase significantly. Although we have limited our use of the word precursor to specifically mean the gamma-ray flashes occurring before the main event (see Tab. 3), it is hoped that such terminology will prove ambiguous in the future. Tables 1 and 2 of [177] provide a summary of next-generation telescopes that could coincidentally detect signals associated with neutrino bursts originating from a binary merger. One intriguing possibility comes from the planned POLAR-2 detector, which should able to measure polarization and similar effects related to the propagation of gamma-rays [a survey of other next-generation gamma-ray facilities are described in Section 2 of 202]. This could be used to probe the local magnetic field strength in the region where the rays were produced. This would allow one to test, for instance, the extent to which non-thermality is tied to magnetars in mergers (see also Sec. IV.4).